OpenStudy (anonymous):

What is the value of y so that the ordered pair (5, y) is a solution to the equation –5x + y = –40? A. –65 B. 9 C. –15 D. 7

OpenStudy (anonymous):

Substitute 5 for x in the equation and solve for y -5(5)+y = -40 -25 + y = -40 y = -40 + 25 y = -15 so the ordered pair is (5,-15)